OpenStudy (anonymous):

Verify [(cot alpha/sec alpha- tan alpha)-(cos alpha/sec alpha +tan alpha)]= sin alpha+csc alpha

zepdrix (zepdrix):

Hmm let's see if we can get this formatted correctly,\[\Large\rm \frac{\cot \alpha}{\sec \alpha- \tan \alpha}-\frac{\cos \alpha}{\sec \alpha+\tan \alpha}=\sin \alpha+\csc \alpha\]Does that look close maybe?

OpenStudy (anonymous):

Yes, it does/

zepdrix (zepdrix):

These identities can be tough sometimes. There are so many ways to approach them. May first instinct is to recognize that the denominators are conjugates of one another, so maybe we can get a common denominator. Hmm thinking whether that's the way we want to go or not...

OpenStudy (anonymous):

So, multiply the LHS fractions by their conjugate?

zepdrix (zepdrix):

No that path seems bad. Kinda makes the problem messier. Let's try this instead,\[\Large\rm \frac{\cot \alpha}{\color{orangered}{\sec \alpha- \tan \alpha}}-\frac{\cos \alpha}{\color{orangered}{\sec \alpha+ \tan \alpha}}=\sin \alpha+\csc \alpha\]Let's write these orange parts in terms of sines and cosines.

zepdrix (zepdrix):

My first instinct was wrong hehe. So anyway, sines and cosines gives us,\[\Large\rm \frac{\cot \alpha}{\color{orangered}{\frac{1}{\cos \alpha}-\frac{\sin \alpha}{\cos \alpha}}}-\frac{\cos \alpha}{\color{orangered}{\sec \alpha+ \tan \alpha}}=\sin \alpha+\csc \alpha\]No no no no no. Ok ok the conjugate thing would have been a good idea actually :P

zepdrix (zepdrix):

Ok sorry sorry I keep flipping back and forth lol.

OpenStudy (anonymous):

No, you're fine. I really appreciate the help!

OpenStudy (anonymous):

Would you also write the second denominator in terms of sines and cosines?

zepdrix (zepdrix):

Ah yes ok fine we'll stick with that path ^^ hehe. Yes the other denominator as well.

OpenStudy (anonymous):

Whichever path you prefer is fine.

zepdrix (zepdrix):

\[\Large\rm \frac{\cot \alpha}{\color{orangered}{\frac{1}{\cos \alpha}-\frac{\sin \alpha}{\cos \alpha}}}-\frac{\cos \alpha}{\color{orangered}{\frac{1}{\cos \alpha}+\frac{\sin \alpha}{\cos \alpha}}}=\sin \alpha+\csc \alpha\] \[\Large\rm \frac{\cot \alpha}{\color{orangered}{\frac{1-\sin \alpha}{\cos \alpha}}}-\frac{\cos \alpha}{\color{orangered}{\frac{1+\sin \alpha}{\cos \alpha}}}=\sin \alpha+\csc \alpha\]See how both terms have a cosine in the bottom? So we can combine those fractions.

zepdrix (zepdrix):

Oh yes yes this path is going to be bad... oh boy.. let's not keep going down this :C My bad my bad.

zepdrix (zepdrix):

Back tot he cojugates.

OpenStudy (anonymous):

It's fine. We can go throught the conjugate path.

zepdrix (zepdrix):

\[\large\rm \frac{\cot \alpha\color{royalblue}{(\sec \alpha+\tan \alpha)}}{(\sec \alpha- \tan \alpha)\color{royalblue}{(\sec \alpha+\tan \alpha)}}-\frac{\cos \alpha\color{#CC0033}{(\sec \alpha-\tan \alpha)}}{(\sec \alpha+\tan \alpha)\color{#CC0033}{(\sec \alpha-\tan \alpha)}}\]

zepdrix (zepdrix):

So our denominators become \(\Large\rm \sec^2\alpha-\tan^2\alpha\)

zepdrix (zepdrix):

Let's look back at our Pythagorean Identity and see if we can do anything with that.

zepdrix (zepdrix):

\[\Large\rm 1+\tan^2x=\sec^2x\]Can we use this to simplify our denominator maybe?

OpenStudy (anonymous):

Okay, that works.

zepdrix (zepdrix):

So what does our denominator simplify to? ;)

zepdrix (zepdrix):

Comeon lyss! You can do this step! I know you can! :3

OpenStudy (anonymous):

Wouldn't that equal positive one? At least with sec^2-tan^2.

zepdrix (zepdrix):

Yes good.

zepdrix (zepdrix):

\[\Large\rm \frac{\cot \alpha(\sec \alpha+\tan \alpha)}{1}-\frac{\cos \alpha(\sec \alpha-\tan \alpha)}{1}\]

zepdrix (zepdrix):

So that deals with the fractions entirely, which is nice.\[\Large\rm \cot \alpha(\sec \alpha+\tan \alpha)-\cos(\sec \alpha-\tan \alpha)\]

zepdrix (zepdrix):

So I guess we need to expand these out. It might be a good idea to convert to sines and cosines at this point. Or maybe after the distributing, I dunno.

zepdrix (zepdrix):

\[\Large\rm \cot \alpha \sec \alpha + \cot \alpha \tan \alpha- \cos \alpha \sec \alpha+ \cos \alpha \tan \alpha\]Oh boy :d

zepdrix (zepdrix):

Look at these two middle terms a sec, \[\Large\rm \cot \alpha \sec \alpha \color{orangered}{+ \cot \alpha \tan \alpha- \cos \alpha \sec \alpha}+ \cos \alpha \tan \alpha\]The first one should simply down really nicely. Remember how to convert cotangent and tangent to sines and cosines?

OpenStudy (anonymous):

Yes, \[(\cos \alpha/ \sin \alpha)(\sin \alpha/\cos \alpha)\]

zepdrix (zepdrix):

Mmm k, and does that simplfy? :)

OpenStudy (anonymous):

\[(1/\sin \alpha)(\sin \alpha0\]

OpenStudy (anonymous):

the zero wasn't suppose to be in there. Sorry.

zepdrix (zepdrix):

Ok good. It should simplify further though. The cosines cancelled out. The sines can cancel out as well, yes? (One is in top-ish, the other in the bottom).

OpenStudy (anonymous):

Yeah. I was wondering if they also cancelled out.

OpenStudy (anonymous):

So, all we're left with is cot alpha*sec alpha+cos alpha*tan alpha?

zepdrix (zepdrix):

Ok so you were able to work ahead a little bit? Each orange term gives you 1, they add to 0. Good good good.

zepdrix (zepdrix):

Yup! Just a tiny bit of simplification to wrap this one up!

OpenStudy (anonymous):

Alright, I believe I have it from here. Thank you!!

zepdrix (zepdrix):

Yay team!